Calculate the force exerted by mars on its moon Phobos and the force exerted by Phobos on mars . The mass of mars is 6.4 x 1023 kg and the mass of Phobos is 1.06 x 1016kg. The distance between bars and the Phobos centers of mass is 9.38 x 106 m


1
Expert's answer
2021-12-21T12:16:04-0500

Let's find the magnitudes of the force exerted by Mars on its moon Phobos and the force exerted by the Phobos on the Mars:


"F_{MP}=F_{PM}=\dfrac{GM_MM_P}{d^2},""F_{MP}=F_{PM}=\dfrac{6.67\times10^{-11}\ \dfrac{N\times m^2}{kg^2}\times6.4\times10^{23}\ kg\times1.06\times10^{16}\ kg}{(9.38\times10^6\ m)^2},""F_{MP}=F_{PM}=5.14\times10^{15}\ N."

The Phobos will exert an equal and opposite force on the Mars.